X has a normal distribution with the given mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 42, σ = 20, find

P(35 ≤ X ≤ 50)
Answer #1

\mu = 42

\sigma = 20

To find P(35 \leq X \leq 50):

Case1: For X from 35 to mid value:
Z + (35 - 42)/20 = - 0.35

Table of Area Under Standard Normal Curve gives area = 0.1368

Case2: For X from mid value to 50:
Z = (50 - 42)/20 = 0.40

Table gives area = 0.1554

So,

P(35 \leq X \leq 50) = 0.1368 + 0.1554 = 0.2922

So,

Answer is:

0.2922